OpenStudy (anonymous):

Need to simplify square root of s/s^2. Answer is s-3/2 but don't know how they got it! Can someone help?

OpenStudy (anonymous):

you mean sqrt(1/s) = s - (3/2) ?

OpenStudy (anonymous):

\[\sqrt{s}\div s ^{2}\]

OpenStudy (nowhereman):

Assuming s is in the correct domain: \[\frac{\sqrt{s}}{s^2} = s^{\frac{1}{2}}\cdot s^{-2} = s^{-\frac{3}{2}}\]

OpenStudy (mathteacher1729):

\(\sqrt{x}=x^{1/2}\) Also \(\frac{x^a}{x^b}=x^{a-b}\) So putting those together... \[\frac{\sqrt{x}}{x^2}=\frac{x^{1/2}}{x^2}=x^{1/2 - 2}=x^{-3/2}\] Hope this helps. :)

OpenStudy (anonymous):

I thought you subtracted the exponents when they were in the "divided by" form! No?

OpenStudy (nowhereman):

You do, I just wrote it in a different form.

OpenStudy (anonymous):

Ok! I see now! I do! Thanks a bunch guys!!! thanks so much!